i am already in a after school toutoring class called compass cause i failed last semester can you please help me with this for free i really need help and i dont have money
my question is
X+5=3x+1

Answer by Simnepi(216)

we don't need money, we just like to help!
to solve any type of equation with an unknown on both sides of the equals sign....
look for which side of the equals sign has the least unknowns
so in the equation
x+5=3x+1
the right hand side (of the equals sign) has 3x and the left hand side has just one x
so we take x from both sides to eliminate the x from the left hand side
giving
x+5-x=3x+1-x
simplifying gives
5=2x+1
Now we have to eliminate the number from the side with the unknown on!
So we have to take 1 from both sides to get rid of the +1 from the right hand side
giving
5-1=2x+1-1
simplifying gives
4=2x
dividing by 2 gives us
2=x or x=2
I hope this helps you. I know algebra can seem hard but keep trying, you'll get there in the end!